OpenStudy (nsh4267):

Solve the equation. ln x=3-ln(x+2)

OpenStudy (nsh4267):

Help!!!

OpenStudy (anonymous):

\[\ln x=3-\ln (x+2)\]\[\ln x+\ln (x+2)=3\]\[\ln (x)(x+2)=3\]\[e(\ln (x)(x+2))=e ^{3}\]\[(x)(x+2)=e ^{3}\]\[x ^{2}+2x=e ^{3}\]\[x ^{2}+2x-e ^{3}=0\]\[x=-1\pm \sqrt{1+e ^{3}}\]